# The local moduli of Sasaki-Einstein $\mathbb{Q}$-homology 7-spheres and invertible polynomials

### v1

@@ 3.2. $\mathbb{Q}$-homology 7-spheres: Thom-Sebastiani sums of invertible polynomials
From the Johnson and Kollar’s list [19] of well-formed anticanonically embedded Fano 3-folds of index 1, one can deduce that the weight vectors associated to these families of polynomials have the form:
$\textbf{w}=(w_{0},w_{1},w_{2},w_{3},w_{4})=(m_{3}v_{0},m_{3}v_{1},m_{2}v_{2},m_{2}v_{3},m_{2}v_{4}),$ (3.2.1)
where $\gcd(m_{2},m_{3})=1$ and, from well-formedness, $\gcd(v_{0},v_{1})=1$ and $\gcd(v_{i},v_{j})=1$ for $i\neq j$ with $i,j\in\{2,3,4\}$. Moreover, it is known that when a weight vector w as in (3.2.1) admits an invertible polynomial $f$ of degree $d=m_{2}m_{3}$ and index $I=|\textbf{w}|-d=1$, then the topology of its link $L_{f}$ correspond to a $\mathbb{Q}$-homology sphere, whose torsion can be calculated from the weights $w_{i}$’s.
Before we find the generator of the space of deformations of the orbifold $Z_{f}=\{f=0\}/\mathbb{C}^{*}(\textbf{w})$, we will prove some technical lemmas on the different types of polynomial described above.
@@
    The equalities above imply $v_{0}\mid m_{2}$ and $v_{1}\mid m_{2}$. Since $w_{i}=m_{2}v_{i}$ for $i=2,3,4$, we have $v_{0}\mid\gcd(w_{0},w_{1},w_{2},w_{3},w_{4})$. Finally, since w is well-formed, we get $v_{0}=1$. $v_{1}=1$ is obtained in a similar way.
@@
Let $f$ be an invertible polynomial of type $I$, $II$ or $III$, where its associated weight vector $\bf w$ is described as in (3.2.1) and its degree satisfies $d=m_{2}m_{3}$. Then the following hold:
@@
For any invertible polynomial $f$ of type I, II or III, whose associated weight vector $\bf w$ is defined as in (3.2.1) and with degree is $d=m_{2}m_{3}$, we have that $f$ contains no block of the form $z_{i}^{2}+z_{j}^{2}$ (recall $\gcd(m_{2},m_{3})=1$). This implies that $\dim\mathfrak{Aut}(Z_{f})=0$, see [7]. Thus, for the hypersurface $\{f=0\}\subset\mathbb{C}^{5}$, the complex dimension of the orbifold $Z_{f}\subset\mathbb{P}(\bf w)$ is given by the formula
$\dim H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))-\sum_{i}\dim H^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(w_{i})).$ (3.2.2)
@@
Finally, as $x_{2}>0$ and $x_{3},x_{4}\in\mathbb{Z}_{0}^{+}$, the equality above implies that $m_{2}=0$, which is not possible. For the converse, the process is similar to that in the previous case. ∎
@@
    $-\dfrac{1}{v_{0}}\leqslant k\leqslant\dfrac{a_{0}}{v_{1}}$ (3.2.9)
@@
    Using again that $\bf w$ does not admit polynomial of type II, we have $v_{1}\neq 1$, which implies that $\dfrac{1}{v_{1}}<1$. In addition, since the link $L_{f}$ is a $\mathbb{Q}$-homology sphere, we have that $\dfrac{m_{2}}{v_{0}v_{1}}-\dfrac{1}{v_{0}}-\dfrac{1}{v_{1}}$ is a integer. Therefore, in (3.2.9) we have:
@@
    Notice that if $f$ is a polynomial of type II such that its associated weight vector $\bf w$ admits a polynomial of type I, then we can work as in a) of lemma above. Therefore, the equation (3.2.4) has $m_{2}+1$ solutions.
@@
    Finally, if $f$ is a polynomial of type III whose associated weight vector admits polynomial of type I and II, then the number of solutions of (3.2.4) is obtained as in a) of the previous lemma.
@@
In addition, as the link $L_{f}$ is a $\mathbb{Q}$-homology sphere, it verifies Equality (2.1.4): $m_{3}=a_{2}a_{3}a_{4}+1$. Thus, we can write $v_{2}$, $v_{3}$ and $v_{4}$ as
@@
$-(a_{2}-1)\leqslant sv_{4}\Rightarrow-\dfrac{a_{2}-1}{v_{4}}\leqslant s.$
@@
Remark 3.2.2.
From the previous two lemmas we have that for $f$ be an invertible polynomial with associated weight vector $\textbf{w}=(w_{0},w_{1},w_{2},w_{3},w_{4})$ satisfying (3.2.1) and with degree $d=m_{2}m_{3}$.
  • If $f$ is a polynomial of type I, then the set of generators of $H^{0}(\mathbb{P}(\textbf{w},\mathcal{O}(d))$ is given by the following monomials of degree $d$:
@@
  • If $f$ is a polynomial of type II and its associated weight vector $\bf w$ does not admit polynomial of type I, then the set of generators of $H^{0}(\mathbb{P}(\textbf{w},\mathcal{O}(d))$ is given by the following monomials of degree $d:$
    $\left\{z_{0}^{m_{2}-kv_{1}}z_{1}^{k},z_{4}z_{2}^{a_{2}},z_{2}z_{3}^{a_{3}},z_{3}z_{4}^{a_{4}},\mbox{ where }0\leqslant k\leqslant\left\lfloor\dfrac{m_{2}}{v_{1}}\right\rfloor+1\right\}.$
  • If $f$ is a polynomial of type III and its associated weight vector $\bf w$ does not admit polynomial of type II,, then the set of generators of $H^{0}(\mathbb{P}(\textbf{w},\mathcal{O}(d))$ is given by the following monomials of degree $d:$
    $\left\{z_{0}^{a_{0}-kv_{1}}z_{1}^{1+kv_{0}},z_{4}z_{2}^{a_{2}},z_{2}z_{3}^{a_{3}},z_{3}z_{4}^{a_{4}},\mbox{ where }0\leqslant k\leqslant\left\lfloor\dfrac{m_{2}}{v_{0}v_{1}}\right\rfloor+1\right\}.$
Now we will compute $\dim H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(w_{i}))$ for each $w_{i}$, that is, we will find all the solutions of the following Diophantine equation
@@
    where $\hat{y}_{0}=y_{0}+1\geqslant 1$ and $\hat{y}_{1}=y_{1}+a_{1}-1\geqslant a_{1}-1$. By the of solutions given above, we can write $\hat{y}_{0}=m_{2}-tv_{1}$ and $\hat{y}_{1}=t$. Thus, we obtain
@@
Since $a_{3}>1$, we have only one solution $(0,0,1,a_{3},0)$. From $\hat{x}_{3}=a_{3}$, we obtain $x_{3}=0$. Then the solution of Equation (3.2.30) is $(0,0,1,0,0)$. Similarly we obtain the solutions $(0,0,0,1,0)$ and (0,0,,0,1)(0,0,,0,1)( 0 , 0 , , 0 , 1 ) for $w_{3}$ and $w_{4}$ respectively. The last statement of the lemma follows for similar arguments as the ones given in the previous lemmas. ∎
In the same vein as Remark 3.1.1 we give the next remark on the moduli of the orbifold associated to the polynomials considered in this subsection.
Remark 3.2.4.
The group $G(\mathbf{w})$ of complex automorphisms of the graded ring $S(\mathbf{w})$ with $\operatorname{Proj}(S(\mathbf{w}))$ can be defined on generators for polynomials of type I, II and III thanks to Lemmas 3.2.6 and 3.2.7. Notice that the generating set of $H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))$ in all these cases is given by quasi-smooth monomials, so we can describe the local moduli of the corresponding orbifold, a $\mathbb{Q}$-homology projective 3-space, as before. Indeed, let $\alpha_{i},\beta_{1}\in\mathbb{C}^{*}$ and $\mathbb{A}\in GL(2,\mathbb{C}),$ then we have
@@
    and the moduli of the orbifold $Z_{f}$ can be expressed by the quotient
@@
    and the moduli of the orbifold $Z_{f}$ can be expressed by the quotient
    $Span\left\{z_{0}^{m_{2}-kv_{1}}z_{1}^{k},z_{4}z_{2}^{a_{2}},z_{2}z_{3}^{a_{3}},z_{3}z_{4}^{a_{4}},\mbox{ where }0\leqslant k\leqslant\left\lfloor\dfrac{m_{2}}{v_{1}}\right\rfloor+1\right\}/G(\mathbf{w}).$
@@
    and the moduli of the orbifold $Z_{f}$ can be expressed by the quotient
    $Span\left\{z_{0}^{a_{0}-kv_{1}}z_{1}^{1+kv_{0}},z_{4}z_{2}^{a_{2}},z_{2}z_{3}^{a_{3}},z_{3}z_{4}^{a_{4}},\mbox{ where }0\leqslant k\leqslant\left\lfloor\dfrac{m_{2}}{v_{0}v_{1}}\right\rfloor+1\right\}/G(\mathbf{w}).$
For the local moduli of Sasaki-Einstein structures for $\mathbb{Q}$-homology spheres we can say that this is non-trivial, more precisely we have:
Theorem 3.2.1.
@@
    If $f$ is of type II and its associated weight vector $\bf w$ does not admit polynomial of type I, then $\mu=\dfrac{m_{2}}{v_{1}}-\dfrac{1}{v_{1}}-2$
@@
    If $f$ is of type III and its associated weight vector $\bf w$ does not admit polynomial of type II, then $\mu=\dfrac{m_{2}}{v_{0}v_{1}}-\dfrac{1}{v_{0}}-\dfrac{1}{v_{1}}-1$.
@@
Moreover, the generators of the spaces $H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))$ and $H^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(w_{i}))$ for all $i=0,\ldots,4$ are given in the previous lemmas of this subsection. Additionally, if $f$ belongs to one of the 236 $\mathbb{Q}$-homology spheres admitting Sasaki-Einstein metrics found in [5] and [13], then the real dimension of the local moduli of Sasaki-Einstein metrics of $\mathbb{Q}$-homology 7-spheres at $L_{f}$ equals $2\left[\dfrac{m_{2}}{v_{0}v_{1}}-\dfrac{1}{v_{0}}-\dfrac{1}{v_{1}}-1\right]$.
@@
    If $f$ is of type I, from Lemmas 3.2.4 and 3.2.5 we know $\dim H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))=m_{2}+4$. Also, from Lemmas 3.2.6 and 3.2.7, we have
    $\sum_{i=0}^{4}\dim H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(w_{i}))=7.$
@@
    $\mu=\dim H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))-\sum_{i}\dim H^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(w_{i}))=m_{2}-3.$
@@
    If $f$ is of type II with associated weight vector $\bf w$ not admitting a polynomial of type I, from Lemmas 3.2.4 and 3.2.5 we know $\dim H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))=\dfrac{m_{2}}{v_{1}}-\dfrac{1}{v_{1}}+4.$ On the other hand, Lemmas 3.2.6 and 3.2.7, establish the equality
    $\sum_{i=0}^{4}\dim H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(w_{i}))=6.$
@@
    $\mu=\dim H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))-\sum_{i}\dim H^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(w_{i}))=\dfrac{m_{2}}{v_{1}}-\dfrac{1}{v_{1}}-2.$
@@
    $\dim H^{0}(\mathbb{P}(\textbf{w}),d)=\dfrac{m_{2}}{v_{0}v_{1}}-\dfrac{1}{v_{0}}-\dfrac{1}{v_{1}}+4.$
@@
    $\sum_{i=0}^{4}\dim H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(w_{i}))=5.$
@@
    $\mu=\dim H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))-\sum_{i}\dim H^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(w_{i}))=\dfrac{m_{2}}{v_{0}v_{1}}-\dfrac{1}{v_{0}}-\dfrac{1}{v_{1}}-1.$
Aso we notice that for any of the cases mentioned above, $\mu$ is equivalent to
$\dim H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))-\sum_{i}\dim H^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(w_{i}))=\dfrac{m_{2}}{v_{0}v_{1}}-\dfrac{1}{v_{0}}-\dfrac{1}{v_{1}}-1.$

### v2

@@ 3.2. Rational homology 7-spheres: Thom-Sebastiani sums of invertible polynomials
Type I (Fermat-Cycle): $f=z_{0}^{a_{0}}+z_{1}^{a_{1}}+z_{4}z_{2}^{a_{2}}+z_{2}z_{3}^{a_{3}}+z_{3}z_{4}^{a_{4}}$
Type II (Chain-Cycle): $f=z_{0}^{a_{0}}+z_{0}z_{1}^{a_{1}}+z_{4}z_{2}^{a_{2}}+z_{2}z_{3}^{a_{3}}+z_{3}z_{4}^{a_{4}}$
Type III (Cycle-Cycle): $f=z_{1}z_{0}^{a_{0}}+z_{0}z_{1}^{a_{1}}+z_{4}z_{2}^{a_{2}}+z_{2}z_{3}^{a_{3}}+z_{3}z_{4}^{a_{4}}.$
Let us assumed that the weight vectors associated to these families of polynomials have the form:
$\mathbf{w}=\left(w_{0},w_{1},w_{2},w_{3},w_{4}\right)=\left(m_{3}v_{0},m_{3}v_{1},m_{2}v_{2},m_{2}v_{3},m_{2}v_{4}\right)$ (3.2.1)
where $\gcd(m_{2},m_{3})=1$ and degree $d=m_{2}m_{3}.$ We will impose the conditions $\gcd(v_{0},v_{1})=1$ and $\gcd(v_{i},v_{j})=1$ for $i\neq j$ with $i,j\in\{2,3,4\},$ which in particular implies that the weighted hypersurface $Z_{f}\subset\mathbb{P}(\bf{w})$ is well-formed.
Before we find the generators of the space of deformations of the orbifold $Z_{f}$, we will prove some technical lemmas on the different types of polynomials described above.
@@
    The equalities above imply $v_{0}\mid m_{2}$ and $v_{1}\mid m_{2}$. Since $w_{i}=m_{2}v_{i}$ for $i=2,3,4$, we have $v_{0}\mid\gcd(w_{0},w_{2},w_{3},w_{4})$. Finally, since w is well-formed, we have $v_{0}=1$. A similar argument shows that $v_{1}=1$.
@@
Let $f$ be an invertible polynomial of type II, IIII or $III$, where its associated weight vector $\bf w$ is described as in (3.2.1) and its degree satisfies $d=m_{2}m_{3}$. Then the following hold:
@@
For any invertible polynomial $f$ of type I, II or III, whose associated weight vector $\bf w$ is defined as in (3.2.1) and with degree $d=m_{2}m_{3}$, we have that $f$ contains no block of the form $z_{i}^{2}+z_{j}^{2}$ (recall $\gcd(m_{2},m_{3})=1$). This implies that $\dim\mathfrak{Aut}(Z_{f})=0$, see [7]. Thus, for the hypersurface $\{f=0\}\subset\mathbb{C}^{5}$, the complex dimension of the orbifold $Z_{f}\subset\mathbb{P}(\bf w)$ is given by the formula
$\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))-\sum_{i}\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(w_{i})).$ (3.2.2)
@@
Finally, as $x_{2}>0$ and $x_{3},x_{4}\in\mathbb{Z}_{0}^{+}$, the equality above implies that $m_{2}=0$, which is not possible. For the converse, the process is similar to the process in the previous argument. ∎
@@
    $\dfrac{1}{v_{0}}\leqslant k\leqslant\dfrac{a_{0}}{v_{1}}$ (3.2.9)
@@
    Using again that $\bf w$ does not admit polynomial of type II, we have $v_{1}\neq 1$, which implies that $\dfrac{1}{v_{1}}<1$. In addition, since $\alpha(\textbf{w})=\dfrac{m_{2}}{v_{0}v_{1}}-\dfrac{1}{v_{0}}-\dfrac{1}{v_{1}}$ is a integer, from (3.2.9) we obtain
@@
    Notice that if $f$ is a polynomial of type II such that its associated weight vector $\bf w$ admits a polynomial of type I, then we can work as in a) of the lemma above. Therefore, the equation (3.2.4) has $m_{2}+1$ solutions.
@@
    Finally, if $f$ is a polynomial of type III whose associated weight vector admits a polynomial of type I and II, then the number of solutions of (3.2.4) is obtained as in a) of the previous lemma.
@@
In addition, as the link $L_{f}$ is a rational homology sphere, it verifies Equality (2.1.4): $m_{3}=a_{2}a_{3}a_{4}+1$. Thus, we can write $v_{2}$, $v_{3}$ and $v_{4}$ as
@@
$(a_{2}-1)\leqslant sv_{4}\Rightarrow-\dfrac{a_{2}-1}{v_{4}}\leqslant s.$
@@
Thus, for $f$ be an invertible polynomial with associated weight vector $\textbf{w}=(w_{0},w_{1},w_{2},w_{3},w_{4})$ satisfying (3.2.1) and with degree $d=m_{2}m_{3}$ we have
  • If $f$ is a polynomial of type I, then the set of generators of $H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))$ is given by the following monomials of degree $d$:
@@
  • If $f$ is a polynomial of type II and its associated weight vector $\bf w$ does not admit polynomial of type I, then the set of generators of $H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))$ is given by the following monomials of degree $d:$
    $\left\{z_{0}^{m_{2}-kv_{1}}z_{1}^{k},z_{4}z_{2}^{a_{2}},z_{2}z_{3}^{a_{3}},z_{3}z_{4}^{a_{4}},\mbox{ where }0\leqslant k\leqslant\left\lfloor\dfrac{m_{2}}{v_{1}}\right\rfloor\right\}.$
  • If $f$ is a polynomial of type III and its associated weight vector $\bf w$ does not admit polynomial of type II, then the set of generators of $H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))$ is given by the following monomials of degree $d:$
    $\left\{z_{0}^{a_{0}-kv_{1}}z_{1}^{1+kv_{0}},z_{4}z_{2}^{a_{2}},z_{2}z_{3}^{a_{3}},z_{3}z_{4}^{a_{4}},\mbox{ where }0\leqslant k\leqslant\left\lfloor\dfrac{m_{2}}{v_{0}v_{1}}\right\rfloor\right\}.$
Now we will compute $\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(w_{i}))$ for each $w_{i}$, that is, we will find all the solutions of the following Diophantine equation
@@
    where $\hat{y}_{0}=y_{0}+1\geqslant 1$ and $\hat{y}_{1}=y_{1}+a_{1}-1\geqslant a_{1}-1$. By the solutions given above, we can write $\hat{y}_{0}=m_{2}-tv_{1}$ and $\hat{y}_{1}=t$. Thus, we obtain
@@
Since $a_{3}>1$, we have only one solution $(0,0,1,a_{3},0)$. From $\hat{x}_{3}=a_{3}$, we obtain $x_{3}=0$. Then the solution of Equation (3.2.30) is $(0,0,1,0,0)$. Similarly we obtain the solutions $(0,0,0,1,0)$ and $(0,0,0,0,1)$ for $w_{3}$ and $w_{4}$ respectively. The last statement of the lemma follows for similar arguments as the ones given in the previous lemmas. ∎
In the same vein as Proposition 3.1.1 we give the following result.
Proposition 3.2.1.
The group $G(\mathbf{w})$ of complex automorphisms of the graded ring $S(\mathbf{w})$ with $\operatorname{Proj}(S(\mathbf{w}))$ can be defined on generators for polynomials of type I, II and III thanks to Lemmas 3.2.6 and 3.2.7 and we can describe the moduli of the corresponding orbifold as before. Indeed, let $\alpha_{i},\beta_{1}\in\mathbb{C}^{*}$ and $\mathbb{A}\in GL(2,\mathbb{C}),$ then we have
@@
    and the moduli of the orbifold $Z_{f}$ is included in the quotient
@@
    and the moduli of the orbifold $Z_{f}$ is included in the quotient
    $Span\left\{z_{0}^{m_{2}-kv_{1}}z_{1}^{k},z_{4}z_{2}^{a_{2}},z_{2}z_{3}^{a_{3}},z_{3}z_{4}^{a_{4}},\mbox{ where }0\leqslant k\leqslant\left\lfloor\dfrac{m_{2}}{v_{1}}\right\rfloor\right\}/G(\mathbf{w}).$
@@
    and the moduli of the orbifold $Z_{f}$ is included in the quotient
    $Span\left\{z_{0}^{a_{0}-kv_{1}}z_{1}^{1+kv_{0}},z_{4}z_{2}^{a_{2}},z_{2}z_{3}^{a_{3}},z_{3}z_{4}^{a_{4}},\mbox{ where }0\leqslant k\leqslant\left\lfloor\dfrac{m_{2}}{v_{0}v_{1}}\right\rfloor\right\}/G(\mathbf{w}).$
$\square$
For the local moduli of Sasaki-Einstein structures for rational homology spheres we can say that this is non-trivial, more precisely we have:
Proposition 3.2.2.
@@
    If $f$ is of type II and its associated weight vector $\bf w$ does not admit a polynomial of type I, then $\mu=\dfrac{m_{2}}{v_{1}}-\dfrac{1}{v_{1}}-2$ .
@@
    If $f$ is of type III and its associated weight vector $\bf w$ does not admit a polynomial of type II, then $\mu=\dfrac{m_{2}}{v_{0}v_{1}}-\dfrac{1}{v_{0}}-\dfrac{1}{v_{1}}-1$.
@@
Moreover, the generators of the spaces $H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))$ and $H^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(w_{i}))$ for all $i=0,\ldots,4$ are given in the previous lemmas of this subsection. Additionally, if $f$ belongs to one of the 236 rational homology spheres admitting Sasaki-Einstein metrics found in [5] and [14], then the real dimension of the local moduli of Sasaki-Einstein metrics of rational homology 7-spheres at $L_{f}$ equals $2\left[\dfrac{m_{2}}{v_{0}v_{1}}-\dfrac{1}{v_{0}}-\dfrac{1}{v_{1}}-1\right]$.
@@
    If $f$ is of type I, from Lemmas 3.2.4 and 3.2.5 we know $\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))=m_{2}+4$. Also, from Lemmas 3.2.6 and 3.2.7, we have
    $\sum_{i=0}^{4}\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(w_{i}))=7.$
@@
    $\mu=\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))-\sum_{i}\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(w_{i}))=m_{2}-3.$
@@
    If $f$ is of type II with associated weight vector $\bf w$ not admitting a polynomial of type I, from Lemmas 3.2.4 and 3.2.5 we know $\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))=\dfrac{m_{2}}{v_{1}}-\dfrac{1}{v_{1}}+4.$ On the other hand, Lemmas 3.2.6 and 3.2.7, establish the equality
    $\sum_{i=0}^{4}\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(w_{i}))=6.$
@@
    $\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))-\sum_{i}\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(w_{i}))=\dfrac{m_{2}}{v_{1}}-\dfrac{1}{v_{1}}-2.$
@@
    $\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\textbf{w}),d)=\dfrac{m_{2}}{v_{0}v_{1}}-\dfrac{1}{v_{0}}-\dfrac{1}{v_{1}}+4.$
@@
    $\sum_{i=0}^{4}\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(w_{i}))=5.$
@@
    $\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))-\sum_{i}\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(w_{i}))=\dfrac{m_{2}}{v_{0}v_{1}}-\dfrac{1}{v_{0}}-\dfrac{1}{v_{1}}-1.$
As we notice that for any of the cases mentioned above, $\mu$ is equivalent to
$\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))-\sum_{i}\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(w_{i}))=\dfrac{m_{2}}{v_{0}v_{1}}-\dfrac{1}{v_{0}}-\dfrac{1}{v_{1}}-1.$